OpenStudy (anonymous):

solve (x-1)/5=(3-2x)/4

OpenStudy (anonymous):

x=19/14

OpenStudy (anonymous):

cross multiply \[\frac{x-1}{5}=\frac{3-2x}{4}\] \[4(x-1)=5(3-2x)\] distribute \[4x-4=15-10x\] put x on one side numbers on other \[14x=19\] divide \[x=\frac{19}{14}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

pls click the GOOD ANSWER button if u r satisfied with the answer:))

OpenStudy (anonymous):

i dont see the good answer button where is it located